OpenStudy (anonymous):

If the roots of the equation x^2-10cx - 11d=0 are a,b and those of the x^2-10ax-11b=0 are c and d then find the value of a+b+c+d (a,b,c,d are distinct)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Well first do u know what a distinct number us?

OpenStudy (anonymous):

Sorry is

OpenStudy (anonymous):

i do

OpenStudy (anonymous):

Ok...So, now I ask do u have the slightest clue on how to solve this or no?

OpenStudy (anonymous):

Here hopefully this helps: tell me when to slow down

OpenStudy (anonymous):

Now i might know what im doing lol but not completely sure solve for the values by looking at it, if x^2 - 10cx - 11d = (x + a)(x + b), then a + b = -10c, and a*b = -11d.

OpenStudy (anonymous):

So aftwr you have done that You can do the same with the other equation, x^2 - 10ax - 11d = (x + c)(x + d), c + d = -10a, and c*d = -11d, but with the second problem, if you divide both sides by d, then you get c = -11, so whatever other stuff is there, c = -11.

OpenStudy (anonymous):

Ok u still with meh?

OpenStudy (anonymous):

you misread

OpenStudy (anonymous):

What

OpenStudy (anonymous):

cd=11b... not cd=11d

OpenStudy (anonymous):

Ahhhhh i see thank u for that

ganeshie8 (ganeshie8):

what kind of numbers are a,b,c,d ? are they integers by any chance ?

OpenStudy (anonymous):

I dont think so

OpenStudy (anonymous):

distinct numbers doesn't imply integers? @ganeshie8

OpenStudy (anonymous):

Wait divu see u said d instead of b bc its the secong equation right?

ganeshie8 (ganeshie8):

distinct' real numbers ?

OpenStudy (anonymous):

nope just distinct numbers

OpenStudy (anonymous):

Use Viete's formula for p(x) = ax^2+bx +c the roots r1,r2 satisfy r1 + r2 = -b/a r1*r2 = c/a

OpenStudy (anonymous):

looks like someone did that already

OpenStudy (anonymous):

Yes umm shall i keep going with the equation

OpenStudy (anonymous):

Bc there are two equations

OpenStudy (anonymous):

I believe right?

OpenStudy (anonymous):

I got a+b = 10c , ab=-11d, c+d = 10a, c*d = 11b

OpenStudy (anonymous):

Hey hey dont go ahead of me jk lemme keep going

OpenStudy (anonymous):

came out with a = 5.5(9 +/- sqrt(85) ), b = 110 - a, c = -11, and d = 11-10a, but i am not too sure how accurate that is

OpenStudy (anonymous):

So yes jazy u r correct if u want to do it the not so fun way jk

OpenStudy (anonymous):

your equation has an issue `x^2 - 10cx - 11d = (x + a)(x + b)` should be x^2 - 10cx - 11d = (x - a)(x - b)

OpenStudy (anonymous):

since a,b are the roots of that quadratic

OpenStudy (anonymous):

Ohhh....

OpenStudy (anonymous):

Yes well i dont know about u but im tired from all this math. Ur right jazy i think idk maybe im wrong

OpenStudy (anonymous):

x^2 - 10cx - 11d = (x - a)(x - b) x^2 - 10cx -11d = x^2 -(a+b)x + ab equating coefficients -10c = -(a+b) 10c = a+b

OpenStudy (anonymous):

Maybe i did my math wrong

OpenStudy (anonymous):

So am i correct

OpenStudy (anonymous):

what did you get

OpenStudy (anonymous):

Here i got a new answer maybe i did my cals again: x2-10ax-11d = (x-c)(x-d) -10a = -c-d => c+d = 10a -11d = cd => -11 = c x2-10cx-11d = (x-a)(x-b) -10c = -b-a => a+b = 10c -11d = ab -11(10a-c) = a(10c-a) -110a + 11c = 10ac - a^2 -110a + 11(-11) = 10a(-11) - a^2 a^2 = 121 a = 11, -11

OpenStudy (anonymous):

My brain is fried literally. Who's with me?

OpenStudy (anonymous):

oops i have an error too a+b = 10c , ab=-11d, c+d = 10a, c*d = `-11b`

OpenStudy (anonymous):

Ohhh so am i correct

OpenStudy (anonymous):

you again made that same mistake

OpenStudy (anonymous):

What

OpenStudy (anonymous):

What grade is this math in????

OpenStudy (anonymous):

x2-10ax-11d = (x-c)(x-d) its not the eqn x2-10ax-11b = (x-c)(x-d) this is

OpenStudy (anonymous):

So im correct im pretty sure

OpenStudy (anonymous):

you are getting confused between 'b' and 'd'

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

What grade is this

OpenStudy (anonymous):

it should be simple, but with a trick and i am not getting it

OpenStudy (anonymous):

Yea me either what grade us this lol for the third time jk

OpenStudy (anonymous):

No but seriously

OpenStudy (anonymous):

wolfram has all the solutions if you want

OpenStudy (anonymous):

don't click if you want to keep working on it http://www.wolframalpha.com/input/?i=solve+a%2Bb+%3D+10c+%2C+ab%3D-11d%2C+c%2Bd+%3D+10a%2C+c*d+%3D+-11b

OpenStudy (anonymous):

Ohhh i see i say google it lol

OpenStudy (anonymous):

I should have just done this in the first place duh

OpenStudy (anonymous):

a = -11d/b d = -11b/c substitute a = (-11 (-11b/c) ) / b = 121 / c

OpenStudy (anonymous):

So I guess that problem is solved. Right?

ganeshie8 (ganeshie8):

a,b are roots of x^2-10cx - 11d=0 c,d are roots of x^2-10ax-11b=0 from vieta formulas we have \(a+b=10c\tag{1}\) \(c+d=10a\tag{2}\) \(\implies (a+b)-(c+d) = 10(c-a) \implies b-d = 11(c-a)\tag{3}\) \(a^2-10ac-11d=0\tag{4}\) \(c^2-10ac-11b=0\tag{5}\) \((5)-(4) \) \(c^2-a^2 = 11(b-d) = 11*11(c-a)\implies c+a = 121\tag{6}\) from \((1), (2), (6)\) : \(a+b+c+d = 10(c+a)=10(121)\)

OpenStudy (anonymous):

And where was this when we needes it. @ganeshie8 jk but seriously this could have been over along time ago

OpenStudy (anonymous):

are you in a rush to go somewhere

OpenStudy (anonymous):

So every one did good. Lol.no im just saying my brain hurts

ganeshie8 (ganeshie8):

here i found the solution https://s3.amazonaws.com/upload.screenshot.co/6853f47c79

OpenStudy (anonymous):

@ganeshie8 thanks yep its from a IIT archieve

OpenStudy (anonymous):

Hope we helped @divu.mkr

OpenStudy (anonymous):

yep thanks guys :)

OpenStudy (anonymous):

Or confuesed u lol

OpenStudy (anonymous):

Np

ganeshie8 (ganeshie8):

it seems we just need c and a are distinct

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Yup er doodle0

OpenStudy (anonymous):

i am curious how wolfram got this solution using these 4 equations nice challenge problem http://www.wolframalpha.com/input/?i=solve+a%2Bb+%3D+10*c+%2C+a*b%3D-11d%2C+c%2Bd+%3D+10a%2C+c*d+%3D+-11b%2C+find+a%2Bb%2Bc%2Bd

OpenStudy (anonymous):

Hey hey were done with tjis question no more lol jk but seriously

OpenStudy (anonymous):

:)